OpenStudy (anonymous):

What is the solution to the system? Line 1 x y –2 –4 –1 –1 Line 2 x y 2 3 3 1 ( , )

OpenStudy (anonymous):

@pooja195 @okdutchman7 @Serenity74

OpenStudy (serenity74):

Rise/Run gives you the coordinates. So (Y2- Y1)/ (X2 - X1)

OpenStudy (serenity74):

You are asking for the coordinates right???

OpenStudy (serenity74):

Right???

OpenStudy (anonymous):

lol yea

OpenStudy (anonymous):

Line 1: y= 3x+2 Line 2: y= -2x +7 Solution: 1,5

OpenStudy (anonymous):

right

OpenStudy (serenity74):

Is that the second question???

OpenStudy (anonymous):

no thats what I got

OpenStudy (serenity74):

Ohhhhhhhhhh waiiiitttt nvm I see what you did there you made those coordinates into equations and then solved for the coordinates right?

OpenStudy (anonymous):

yea

OpenStudy (serenity74):

Lemme check if it's right lol gimme a sec.

OpenStudy (anonymous):

am I correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@pooja195

OpenStudy (serenity74):

How did you get 1, 5??

OpenStudy (serenity74):

I got fractions as my answers....

OpenStudy (anonymous):

@Serenity74 is my answer 5,11

OpenStudy (serenity74):

I got fractions.

OpenStudy (serenity74):

You calculated you second equation wrong.

OpenStudy (anonymous):

ok one sec

OpenStudy (serenity74):

No wait! Ilied. I f***ed up.

OpenStudy (anonymous):

lol

OpenStudy (serenity74):

I still got a fraction.

OpenStudy (anonymous):

let me ask one of the smartest people on openstudy not saying your not smart. your pretty smart @Serenity74

OpenStudy (anonymous):

@pooja195 is my sleek-feathered oneaaaaaaaaaaaaaa

OpenStudy (serenity74):

No it's okay. I am not offended lol I know I am not smart XD XD XD XD Don't worry bruh. Lol.

OpenStudy (anonymous):

lol thanks tho @Serenity74

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

is my answer 1,5 or 5,11

OpenStudy (alexandervonhumboldt2):

let me see

OpenStudy (alexandervonhumboldt2):

function rule for the first table is y=3x+2 for second it is -2x+7 equal them 3x+2=-2x+7 5x=5 x=1 substitute 1 instead of x y=3x+2 y=3+2 y=5 so it is (1, 5)

OpenStudy (anonymous):

thakns

OpenStudy (alexandervonhumboldt2):

:)

OpenStudy (anonymous):

thanks*